-- -- UsingBB Introduction Purpose ThisbulletinboardisprovidedbyKenCaseasaservicetoWizardsoftheCoast,incooperation withLisaLowe,PeterAdkisonandKenMcGlothlen.Thesoftwareiswrittenandmaintainedby KenCaseandWilShipley;itwasoriginallywrittenattheUniversityofWashington,andisstill beingusedthere. Asanewuser,youareadvisedtoreadthisdocumentbeforeproceedingfurtherinbb.Command referenceisavailablewithinbbviathehelpcommand. Ethics Theuseofthisbulletinboardisaprivilege--notaright.Usersareremindedthatthecontentof postingsaresubjecttofederalandstatelawsaswellasUniversitypoliciesonsexual,racialand religiousharassment,misuseofUniversityproperty,anddiscriminationbasedonage,veteranstatus,gender,sexualorientation,race,ethnicoriginormaritalstatus--regardlessofthe‘‘historical culture’’oftheroominwhichamessageisposted. Failuretocomplywiththeselawsandpoliciescouldresultindisciplinaryactionsincluding, butnotlimitedto,suspensionorpermanentdismissalfromtheUniversity,civilprosecution, and/orpersonalliability. ComplaintsregardingallegedmisconductshouldbedirectedtoKenMcGlothlenat206/784-2083, ormcglk@cpac.washington.edu. AccessingBB Connectingtobb Toconnecttobbenterthecommand: bb onblegga,or: bb-Sblegga.cac onanyoftheCPACorUCSUNIXmachines. Signingon Whenyouconnecttobbyouwillseethefollowingmessage:

-- -- -2-WelcometoBB. bbid: Ifyouhavealreadyestablishedabbsignon,enteryouridatthisprompt,andyourpasswordat the‘‘password’’promptthatfollows.Ifnot,enter‘‘new’’asyourid.Thiswillproducethefollowing: WelcometoBB. bbid:new Creatinganewbbsignon.Enteranunusedidandpassword. bbid: . . . bbpassword: Atthe‘‘id’’and‘‘password’’prompts,enteranynewbbidandpasswordyouwish.Theidand passwordyouchooseneednotbethesameasyoursystemidorpassword.Tochangeyouridor passwordlater,usethesetnameorpasswordcommandswithinbb. GettingAround Architecture Inbb,messagesarepostedinrooms;eachroomaddressesaspecifictopic.Allroomsareinone oftwogroups:‘‘supported’’and‘‘unsupported.’’ThesupportedgroupcontainsthoseroomssupportedbyUniversityComputingServices.Unsupportedroomsareinitiallyhiddenfromview,and messageswithinthoseroomsexpiremorequicklythanthoseinthesupportedrooms. Rooms:PrioritiesandAccess Bydefault,roomsareaccessiblebyeveryone.However,theycanbeassignedtheclassnoaccess bytheirowner.Iftheyhavethisclass,onlythosespecifiedbytheownercaneverseetheroom. Roomsaresortedbytheirpriority,whichisanumberassignedtoindividualroomsandgroups.A roomassignedahigherprioritycomesbeforearoomwithalowerpriority;roomsassignedthe samepriorityaresortedalphabetically. Aroom’spriorityisdeterminedbythehighestpriorityofthegroupsinwhichthatroomisa member,andanindividualoffsetfortheroom’spriority.Tochangethepriorityofaroom,use theprioritycommand,andspecifywhatyouwouldliketheroom’snewprioritytobe:bbwill computetheappropriateoffsetfortheroom.Forexample, -- -- -3-bb>priorityunix*hints50 setsthepriorityofthe‘‘unix*hints’’roomto50(regardlessofthegroupsofwhichunix*hintshappenstobeamember). Thereisaconfigurablecutoffpriority,whichisinitiallysetto0.Allroomswithprioritiesbelow thecutoffarehidden.Thosewithhigherprioritiesare(ifaccessible)automaticallyvisible.To changeyourcutoffpriority,usethecutoffcommand. Bydefault,onlytheroomsinthesupportedgrouparevisible.Thepriority/groupcommand changesagroup’spriority.Toseetheroomsintheunsupportedgroup,setitsprioritytoa numberaboveyourcutoff,e.g., bb>priority/groupunsupported0 Formoreinformation,seetheinternalhelp. Directory Thedirectorycommandshowsalistofallthevisible,accessiblerooms.Ifmessageshavebeen postedinaroomfollowingthelastmessageyou’vereadinthatroom,itwillbeflaggedwithan asterisk.Thecurrentlyselectedroomisflaggedwith‘‘>’’.Therearevariousqualifiers(typedirectory/<qualifier>)thatmodifydirectory’soutput.Themostcommonlyusedare: /countdisplaysthenumberofmessagespostedineachroomfollowingthelast messagereadinthatroom /fulldisplayseachroom’ssummarylineandowneraswellasthestandardroom name /messageslistsonlythoseroomsinwhichmessageshavebeenpostedsincethelast messageyou’veread /prioritiesdisplaysroompriorities Foracompletelistofqualifiers,seetheinternalhelp. Selectingrooms Whenyouenterbb,youareplacedinthefirstroominyourdirectorywhichcontainsnewmessages(seethediscussionofpriorities,above,forinformationonhowtore-orderyourdirectory). Thecommandlinepromptalwaysshowswhichroomyouarein,themessagenumberyouareat andthetotalnumberofmessagesinaroom.Forexample, bb-general:19/104> tellsyouthatyouareinthe‘‘general’’room,the19thmessageisthecurrentmessage,andthere are104messagesintheroom.Thecurrentmessageisthelatestoneread.

-- -- -4-Toselectadifferentroomtoreadfromand/orpostin,usethe‘‘goto<room-name>’’command. Forexample, bb>gotounix*hints selectsthe‘‘unix*hints’’room.Youcanabbreviateroomnamestoasfewlettersasyouchoose;if whatyouspecifyisambiguous,bbwillselectthefirstmatchingnameinyourdirectory.For example, bb>gotow couldselect‘‘weather,’’‘‘whine,’’‘‘whyever,’’etc.,dependingupontherelativeorderoftheserooms inyourdirectory. Youmayleaveoffthegoto,aslongastheroomnamedoesn’tconflictwithabbcommand.For example,toselectthevms*hintsroom,youmaysimplytype‘‘vms*hints’’(orprobablyevenjust ‘‘vms’’). Readingmessages Thereareseveralcommandsforreadingmessages.Toreadthenextmessageinaroom,usethe nextcommand,orjusthitthe<return>key.Toreadaspecificmessage,type‘‘read<message number>’’orsimply‘‘<messagenumber>’’.(Forexample,typing‘‘23’’willreadthe23rdmessageofthecurrentroom.)Thefirstcommandreadsthefirstmessageinaroom;likewise,last readsthelastone.Toreadthemessagebeforethecurrentone,useback.Allreadsallofthe newmessagesinaroomonerightaftertheother,withoutacommandpromptinbetween. Summary Summaryliststhemessagesinaroom,displayforeachthemessagenumber,theauthorandthe timeofitsposting,anditssubjectline.Tolistonlythosemessagespostedafterthecurrentmessage,usethe/newqualifier. PostingMessages Generalinstructionsforpostingamessage Topostamessage,usethepostcommand.Youwillbepromptedforasubjectline.Trytomake thislinerelevantandinformative:itshouldgiveabriefdescriptionofthepost’ssubjectmatter. Therearetwowaystopostmessages:withaneditor,andwithoutone.Thelatteriseasierinthe shortrun,becauseitdoesnotrequireyoutolearnanyeditors.Theformerismoreusefulinthe longrun;youmaychangemorethings,moreeasily,inyourmessageasyouarewritingit. Ifyoudonotknowhowtouseaneditorandwouldliketolearn,theUNIXmanualpages(accessed viatheUNIXcommandman)offerinstructions,ordocumentationcanbepickedupatacampus consultingoffice. Theeditortobeused,ifany,isdeterminedbytheseteditorcommand,orbythesystemenvironmentvariablesBBEDITORandEDITOR.Formoreinformation,see‘‘helpset.’’Bydefault, bbdoesnotuseaneditor.

-- -- -5-Postingwithoutaneditor Postprintsalineofinstructionstoyourterminalafteryouenteryoursubject: |------Enteryourmessage,followedbyCTRL/Donablankline.------| Trytoendyourlinesbeforetheverticalbarattherightendoftheinstructionline,ormanyterminalswillnotbeabletoreadtheendsofthem.Youcannotchangelinesabovetheoneyouare on,butyoucaneditthelineyouaretypingon,usingthestandardlineeditingkeys(forexample, the<delete>key).ThearrowkeyswillnotworkontheUNIXversion. Whenyou’vefinishedyourmessage,typeControl-D(orControl-Z,onVMS)onablankline.(To typeacontrolcharacter,holddownthe<control>keyandpressthecharacter.)Therearefour relevantcontrolcharactersifyou’renotusinganeditor: Control-Dpostsyourmessage,whentypedonablankline. Control-Lmarksanewpageinyourtext.Whenthemessageappearsonauser’sscreen,it willscrolldowntotheControl-Lcharacterandstop.Hittingthespacebarwill givetheuserthenextpageoftext. Control-Zworksatanypointinbb.Itputsyouatyoursystemprompt;youcangetback ontobbwiththecommandfg.Yourmessagewillstillbeintact,thoughyouwill notbeabletoseewhatyouhavealreadytyped. Control-Cabortsyourmessage:nothingisposted. PostingFiles Topostafiletobb,usethecommand: bb>postsubject<filename where‘‘filename’’isthenameofthefileyouwishtopost.Ifyoudon’tspecifyasubjectonthe commandline,thefirstlineofthefilewillbeusedasthesubject. Deletingmessages Thedeletecommandwillgetridofunwantedmessagesthatyouhaveposted;enter‘‘delete’’ whenthemessageisthecurrentmessage,orenter‘‘delete<messagenumber>’’todeletethe messagespecifiedby<messagenumber>.

-- -- -6-Nicknames Anicknameisashortlineoftextthatappearswithyourbbidinmessagesyoupost.Tosetor changeit,usethesetnickcommand,whichpromptsyouforyournewnickname,andthenrepeats itbacktoyou: bb>setnick nickname?ZoologyStatisticsConsultant nicknamechangedtoZoologyStatisticsConsultant Youmayalsospecifyyournicknamefollowingthesetnickcommandonthecommandline,e.g., ‘‘setnickZoologyStatisticsConsultant’’. Miscellaneous BB’sinitializationfile Whenbbstartsup,itlooksinyourhomedirectoryforafilenamed‘‘.bbrc’’.Ifthisfileexists,it shouldcontainasetofcommandsforbbtoexecuteasitstartsup.Auser’s.bbrcfilemightcontainthefollowingcommands: signonhumboldtocean setpager=less seteditor=emacs users aliasd*ir"directory/messages" aliashi*de"hide!*;go" dir Thefirstlinewillconnecttheuser(humboldt)tobbwithouthimhavingtomanuallytypeinhis idorpassword(‘‘ocean’’). Thesecondlinesetsthepager:inthiscase,less.Apageristheprogramusedtodisplaymost outputfrombb. Thethirdlinedefinestheeditortobeusedwhenpostingmessages.Youreditorandpagerwill defaulttovaluesfoundinyouraccount’senvironmentvariables,ifnotspecifiedwithset.(Seethe internalhelpforsetformoreinformation.) Thefourthline,‘‘users’’,isacommandwhichliststheusersonbb,displayingtheirsystemids, theirnicknames,andtheirbbids. Thenexttwolinesintheinitializationfiledefinealiases: ‘‘aliasd*ir"directory/messages"’’setsupanaliassothat‘‘dir’’invokesthecommand‘‘directory /messages’’,whichdisplaysalistofallthevisible,accessibleroomswithnewmessages. ‘‘aliashi*de"hide!*;go"’’createsanaliassothattyping‘‘hide’’goestothenextroomwithnew messagesafterhidingtheroom.The‘‘!*’’specifiesthatanyargumentstothehidealiasshouldbe -- -- -7-insertedatthatpointinthetranslation. Thelastline,‘‘dir’’,invokesthepreviouslydefinedaliasfordirectory. Formoreinformation,seetheinternalhelpfor‘‘.bbrc’’. Executingshellcommandsfromwithinbb Toexecuteshellcommandswhileinbb,precedeyourshellcommandwitha‘‘!’’.Forexample, bb>!lessblegga scrollsafile(named‘‘blegga’’)inyouraccountthroughlessontoyourscreen,andthenreturnsyou tothebbcommandline.Similarly, bb>!prtguide willprintthefilenamed‘‘guide’’. (Pleasenote:inthecurrentimplementation,shellcommandsareexecutedbytheBourneshell. Oneside-effectofthisisthatcshandtcshaliaseswillnotbetranslated.) FileManipulation ‘‘<’’,‘‘>’’,and‘‘>>’’redirectinputandoutputtoandfromfilesforallbbcommands,justas theydoinUNIXshells.Forexample, bb>all>>bb-support.messages appendsallmessagesinthecurrentroomtoafilenamed‘‘bb-support.messages’’. Formoreinformation,seetheinternalhelpfor‘‘filemanipulation.’’ Wheretogofromhere Ifyouhavequestionsthataren’tansweredinthisdocumention,trylookingthroughtheinternal help,orpostingamessageinthe‘‘bb*questions’’room.Ifyoufindaproblemwiththeprogram, postaboutitinthe‘‘bugs’’room.Suggestionsforimprovingbbcanbepostedinthe‘‘suggestions’’room.

-- -- -8-Listofhelptopics Internalhelpisavailableonthefollowingtopics: Concepts CommandsInitializationHeaderFormatParameters EthicsFileManipulationMessageNumberSpecialEffects Commands changeclasshidequitset createinforeadshow deletelastremovesignon adddirectorylistrenamesource aliasechomailreplysubtract allexitnextresetsummary backfirstpostrespondunalias catchupgotopreservesearchundelete changehelppriorityrunusers ProgramTopics ProgrammingVariablesglobalprompt Macros return Selectorsforletwhile Type‘‘help’’atthebbpromptforanup-to-datelistofavailablehelptopics. -- --